Required information

The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,750 $ 16,650 $ 29,400
Estimated variable manufacturing overhead per machine-hour $ 2.50 $ 3.30
Job P Job Q
Direct materials $ 24,000 $ 13,500
Direct labor cost $ 29,800 $ 11,900
Actual machine-hours used:
Molding 2,800 1,900
Fabrication 1,700 2,000
Total 4,500 3,900

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-9, assume that Sweeten Company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments and Job P included 20 units and Job Q included 30 units. For questions 10-15, assume that the company uses a plantwide predetermined overhead rate with machine-hours as the allocation base.

14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

15. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1

Plantwide overhead rate = (2500*2.5+1500*3.3+29400)/4000 = 10.15 per hour

14) Calculate selling price and unit price

Job P Job Q
Direct material 24000 13500
Direct labor 29800 11900
Overhead applied 4500*10.15 = 45675 3900*10.15 = 39585
Total Cost 99475 64985
Selling price 99475*1.8 = 179055 64985*1.8 = 116973
Unit price 179055/20 = 8953 3899

15) Cost of goods sold = 99475+64985 = 164460